# Vertex F-algebras and their associated Lie algebra

### v1

@@ 3.6 F-delta distributions
3.7 F-residues
Recall that the residue of a formal Laurent series $f(z)=\sum a_{n}z^{n}\in R\llbracket z^{\pm 1}\rrbracket$ is defined as $\operatorname{Res}_{z=0}f(z)dz=a_{-1}.$ We have the identities (see [15, Ch. 13])
$\operatorname{Res}_{z=0}f^{(n)}(z)dz$ $=0,\enskip\forall n\geqslant 1,$ (3.26)
$\operatorname{Res}_{z=0}f^{\prime}(z)g(z)dz$ $=-\operatorname{Res}_{z=0}f(z)g^{\prime}(z)dz,$ (3.27)
$\operatorname{Res}_{w=0}f(h(w))h^{\prime}(w)dw$ $=\operatorname{Res}_{z=0}f(z)dz.$ (3.28)
Here $f,g\in R\llbracket z^{\pm 1}\rrbracket$ are formal Laurent series and $h$ is a (holomorphic) formal power series with $h(0)=0$ and $h^{\prime}(0)\in R^{\times}.$
This terminology can be extended to formal group laws $F(z,w)$ as follows. Recall the invariant $1$-form $\theta_{F}$ from (2.4). Then the F-residue is defined as
$\operatorname{Res}_{z=0}^{F}f(z)dz=\operatorname{Res}_{z=0}f(z)\theta_{F}.$ (3.29)
Assuming the existence of a logarithm,
$\operatorname{Res}_{z=0}^{F}f(z)dz$ $=\operatorname{Res}_{x=0}f(x)dx,$ where $\text{where }f(x)=f(\phi^{-1}(x)).$ (3.30)
From (2.7) we have
$\operatorname{Res}_{z=0}^{F}f(\iota(z))dz=-\operatorname{Res}_{z=0}^{F}f(z)dz.$ (3.31)
From (3.25) combined with (3.21) and (3.22) we find
$\operatorname{Res}_{z=0}^{F}z^{-1}\delta_{F}\left(\frac{w}{z}\right)dz=\operatorname{Res}_{w=0}^{F}z^{-1}\delta_{F}\left(\frac{w}{z}\right)dw=1.$ (3.32)
3.8 F-hyperderivatives
Generalizing Hasse’s number-theoretic notion to formal group laws, define the F-hyperderivative of $f\in R\llparenthesis z\rrparenthesis,$ $\mathcal{D}^{F}_{n}f,$ $n\geqslant 0,$ by expanding the substitution
$i_{z,\underline{w}}f(F(z,w))=\sum\nolimits_{n\geqslant 0}(\mathcal{D}^{F}_{n}f)(z)w^{n}.$ (3.33)
Hence $\mathcal{D}^{F}_{n}f(0)$ picks out the $n$-th coefficient of $f,$ and we think of $\mathcal{D}^{F}_{n}f$ as a substitute for $(n!)^{-1}f^{(n)}(z).$ One easily shows
$(\mathcal{D}_{0}^{F}f)(z)$ $=f(z),$ (3.34)
$(\mathcal{D}_{1}^{F}f)(z)\theta_{F}$ $=f^{\prime}(z)dz,$ (3.35)
$\mathcal{D}_{n}^{F}(\lambda f+g)$ $=\lambda\mathcal{D}_{n}(f)+\mathcal{D}_{n}(g),$ (3.36)
$\mathcal{D}_{n}^{F}(f\cdot g)$ $=\sum\nolimits_{i+j=n}\mathcal{D}_{i}^{F}(f)\cdot\mathcal{D}_{j}^{F}(g),$ (3.37)
$\mathcal{D}_{m}^{F}\mathcal{D}_{n}^{F}(f)$ $=\mathcal{D}_{n}^{F}\mathcal{D}_{m}^{F}(f)=\sum\nolimits_{k\geqslant 0}\binom{k}{m,n}_{\!\!F}\mathcal{D}_{k}^{F}(f).$ (3.38)
For the additive group law, (3.38) implies $(\mathcal{D}_{1}^{F_{a}})^{\circ n}=n!\cdot\mathcal{D}^{F_{a}}_{n},$ which shows the importance of considering higher hyperderivatives in the presence of torsion. Equation (3.38) follows from the associativity of $F.$
@@ 3.8 F-hyperderivatives
$\mathcal{D}_{j}^{F}(z^{m})=\sum_{i\in\mathbb{Z}}\binom{m}{i,j}_{\!\!F}z^{i}.$ (3.39)
From this we see that (3.38) implies
@@
For all $n\geqslant 1$ we have
$\operatorname{Res}^{F}_{z=0}\mathcal{D}_{n}^{F}f(z)dz$ $=0,$ (3.40)
$\operatorname{Res}^{F}_{z=0}\mathcal{D}_{n}^{F}f(z)\cdot g(z)dz$ $=-\sum_{j=1}^{n}\operatorname{Res}_{z=0}^{F}\bigl{[}\mathcal{D}_{n-j}^{F}f(z)\cdot\mathcal{D}_{j}^{F}g(z)\bigr{]}dz.$ (3.41)
@@
$\operatorname{Res}^{F}_{z=0}i_{z,\underline{w}}f(F(z,w))dz=\operatorname{Res}^{F}_{z=0}f(z)dz.$
We will apply the substitution rule (3.28) to the power series $h(z)=F(z,w)-w$ with coefficients in the ring $R\llbracket w\rrbracket.$ This corresponds to viewing $w$ as a fixed constant. Putting $dw=0$ in (2.6) gives $p_{F}(F(z,w))F^{1,0}(z,w)=p_{F}(z).$ Clearly $h(0)=0$ and $h^{\prime}(0)=F^{1,0}(0,w)=p_{F}(w)^{-1}$ is a unit, so the hypotheses for (3.28) are satisfied. Applying this and unravelling the definition gives
$\operatorname{Res}^{F}_{z=0}i_{z,\underline{w}}f(F(z,w))dz$ $=\operatorname{Res}_{z=0}i_{z,\underline{w}}f(F(z,w))p_{F}(z)dz$
$=\operatorname{Res}_{z=0}i_{z,\underline{w}}f(F(z,w))p_{F}(F(z,w))F^{1,0}(z,w)dz$
$=\operatorname{Res}_{z=0}i_{z,\underline{w}}f(h(z)+w)p_{F}(h(z)+w)h^{\prime}(z)dz$
$=\operatorname{Res}_{v=0}i_{v,\underline{w}}f(v+w)p_{F}(v+w)dv$
Let $g(z)=f(z)p_{F}(z)=\sum_{n\in\mathbb{Z}}a_{n}z^{n}\in R\llparenthesis z\rrparenthesis.$ It remains to prove that
$\operatorname{Res}_{v=0}i_{v,\underline{w}}g(v+w)dv=\operatorname{Res}_{z=0}g(z)dz,$
which is a straight-forward computation:
$\operatorname{Res}_{v=0}i_{v,\underline{w}}g(v+w)dv=\operatorname{Res}_{v=0}\sum_{n,k\in\mathbb{Z}}a_{n}\binom{n}{k}v^{n-k}w^{k}dv=a_{-1}$
@@
For all $f\in R\llbracket x_{0}^{\pm 1},x_{1}^{\pm 1},x_{2}^{\pm 1}\rrbracket$ such that the substitutions $f(x_{1}-x_{2},x_{1},x_{2}),$ $f(x_{0},x_{1},x_{1}-x_{0})$ converge in the algebraic sense,
$\operatorname{Res}_{x_{1}=0}\operatorname{Res}_{x_{2}=0}i_{x_{1},x_{2}}f(x_{1}-x_{2},x_{1},x_{2})dx_{2}dx_{1}\\ -\operatorname{Res}_{x_{2}=0}\operatorname{Res}_{x_{1}=0}i_{x_{2},x_{1}}f(x_{1}-x_{2},x_{1},x_{2})dx_{1}dx_{2}\\ =\operatorname{Res}_{x_{1}=0}\operatorname{Res}_{x_{0}=0}i_{x_{1},x_{0}}f(x_{0},x_{1},x_{1}-x_{0})dx_{0}dx_{1}.$ (3.42)
@@
$\operatorname{Res}^{F}_{z_{1}=0}\operatorname{Res}^{F}_{z_{2}=0}i_{z_{1},z_{2}}f(F(z_{1},\iota(z_{2})),z_{1},z_{2})dz_{2}dz_{1}\\ -\operatorname{Res}^{F}_{z_{2}=0}\operatorname{Res}^{F}_{z_{1}=0}i_{z_{2},z_{1}}f(F(z_{1},\iota(z_{2})),z_{1},z_{2})dz_{1}dz_{2}\\ =\operatorname{Res}^{F}_{z_{1}=0}\operatorname{Res}^{F}_{z_{0}=0}i_{z_{1},z_{0}}f(z_{0},z_{1},F(z_{1},\iota(z_{0})))dz_{0}dz_{1}.$ (3.43)
@@
$(-1)^{c-1}\binom{a}{1-c}-(-1)^{a+b-1}\binom{a}{1-b}=(-1)^{a-1}\binom{c}{1-a},$
@@
It suffices to prove (3.43) for the universal formal group law, where we may assume (2.9). In this case, we can use (3.30) to reduce (3.43) to (3.42) for the function $f(x_{0},x_{1},x_{2}),$ where $\phi(z_{i})=x_{i},$ $i=0,1,2.$
4 Vertex F-algebras and their Lie Algebras
@@ 4.1 Axioms
A vertex $F$-algebra over a formal group law $F(z,w)\in R\llbracket z,w\rrbracket$ consists of data $(V,\mathbb{1},\mathcal{D},Y)$ as follows:
  • an $R$-module $V$ of states
  • a vacuum vector $\mathbb{1}\in V$
@@
    $\mathcal{D}(z)\colon V$ $\longrightarrow V\llbracket z\rrbracket,$ $\mathcal{D}(z)a=\sum\nolimits_{n\geqslant 0}\mathcal{D}^{(n)}(a)z^{n}$ (4.1)
@@
    $V\otimes_{R}V$ $\longrightarrow V\llparenthesis z\rrparenthesis,$ $a\otimes b\longmapsto Y(a,z)b$ (4.2)
    $Y(a,z)b$ $=\sum\nolimits_{n\in\mathbb{Z}}a_{(n)}(b)z^{-n-1},$ $a_{(n)}(b)=0$ for $n\gg 0.$ (4.3)
@@
  1. (a)
    Vacuum & creation: $Y(a,z)\mathbb{1}$ is holomorphic for all $a\in V$ and
    $\left.Y(a,z)\mathbb{1}\right|_{z=0}$ $=a,$ (4.4)
    $Y(\mathbb{1},z)$ $=\operatorname{id}_{V}.$ (4.5)
  2. (b)
    $F$-translation covariance: for all $a,b\in V$ we have
    $Y(\mathcal{D}(w)(a),z)b$ $=i_{z,w}Y(a,F(z,w))b,$ (4.6)
    $\mathcal{D}(z)\mathbb{1}$ $=\mathbb{1}.$ (4.7)
    In (4.6) we have used the substitution (3.15). Moreover,
    $\mathcal{D}(z)\circ\mathcal{D}(w)$ $=\mathcal{D}(F(z,w)),$ $\mathcal{D}(0)$ $=\operatorname{id}_{V};$ (4.8)
  3. (c)
    Weak $F$-associativity: for all $a,b,c\in V$ there exists $N\geqslant 0$ with
@@
    using the substitution $v\mapsto F(z,w)$ from (3.13) on the right hand side.
  4. (d)
@@
    $Y(a,z)b=\mathcal{D}(z)\circ Y(b,\iota(z))a.$ (4.10)
Putting $b=\mathbb{1}$ into (4.10) and using (4.5) gives
$Y(a,z)\mathbb{1}=\mathcal{D}(z)a.$ (4.11)
@@
The axioms are slightly redundant. For example, (4.6) implies (4.9) for $b=\mathbb{1}.$ On the other hand, we cannot deduce (4.6) from (4.9) unless we know a priori that both sides of (4.6) are meromorphic in $z,w.$ Hence (4.6) encodes a ‘meromorphicity’ property.
Problem 4.3.

### v2

@@ 3.6 $\boldsymbol{F}$-delta distributions
3.7 $\boldsymbol{F}$-residues
Recall that the residue of a formal Laurent series $f(z)=\sum a_{n}z^{n}\in R\big\llbracket z^{\pm 1}\big\rrbracket$ is defined as $\operatorname{Res}_{z=0}f(z){\rm d}z=a_{-1}$. We have the identities (see [14, Chapter 13])
$\operatorname{Res}_{z=0}f^{(n)}(z){\rm d}z=0,\forall n\geqslant 1,\qquad\operatorname{Res}_{z=0}f^{\prime}(z)g(z){\rm d}z=-\operatorname{Res}_{z=0}f(z)g^{\prime}(z){\rm d}z,$
$\operatorname{Res}_{w=0}f(h(w))h^{\prime}(w){\rm d}w=\operatorname{Res}_{z=0}f(z){\rm d}z.$ (3.20)
Here $f,g\in R\big\llbracket z^{\pm 1}\big\rrbracket$ are formal Laurent series and $h$ is a (holomorphic) formal power series with h(0)=0{h(0)=0} and $h^{\prime}(0)\in R^{\times}$.
This terminology can be extended to formal group laws $F(z,w)$ as follows. Recall the invariant $1$-form $\theta_{F}$ from (2.4). Then the F-residue is defined as $\operatorname{Res}_{z=0}^{F}f(z){\rm d}z=\operatorname{Res}_{z=0}f(z)\theta_{F}$. Assuming the existence of a logarithm,
$\operatorname{Res}_{z=0}^{F}f(z){\rm d}z=\operatorname{Res}_{x=0}f(x){\rm d}x,\qquad\text{where}\quad f(x)=f\bigl(\phi^{-1}(x)\bigr).$ (3.21)
From (2.6), we have
$\operatorname{Res}_{z=0}^{F}f(\iota(z)){\rm d}z=-\operatorname{Res}_{z=0}^{F}f(z){\rm d}z.$ (3.22)
From (3.19) combined with (3.15) and (3.16), we find
$\operatorname{Res}_{z=0}^{F}z^{-1}\delta_{F}\left(\frac{w}{z}\right){\rm d}z=\operatorname{Res}_{w=0}^{F}z^{-1}\delta_{F}\left(\frac{w}{z}\right){\rm d}w=1.$
3.8 $\boldsymbol{F}$-hyperderivatives
Generalizing Hasse’s number-theoretic notion to formal group laws, define the F-hyperderivative of $f\in R(\!(z)\!)$, $\mathcal{S}^{F}_{n}f$, $n\geqslant 0$, by expanding the substitution
$i_{z,\underline{w}}f(F(z,w))=\sum_{n\geqslant 0}\bigl(\mathcal{S}^{F}_{n}f\bigr)(z)w^{n}.$ (3.23)
Hence $\mathcal{S}^{F}_{n}f(0)$ picks out the $n$-th coefficient of $f$, and we think of $\mathcal{S}^{F}_{n}f$ as a substitute for $(n!)^{-1}f^{(n)}(z)$. One easily shows
$\bigl(\mathcal{S}_{0}^{F}f\bigr)(z)=f(z),\qquad\bigl(\mathcal{S}_{1}^{F}f\bigr)(z)\theta_{F}=f^{\prime}(z){\rm d}z,\qquad\mathcal{S}_{n}^{F}(\lambda f+g)=\lambda\mathcal{S}_{n}(f)+\mathcal{S}_{n}(g),$
$\mathcal{S}_{n}^{F}(f\cdot g)=\sum_{i+j=n}\mathcal{S}_{i}^{F}(f)\cdot\mathcal{S}_{j}^{F}(g),$ (3.24)
$\mathcal{S}_{m}^{F}\mathcal{S}_{n}^{F}(f)=\mathcal{S}_{n}^{F}\mathcal{S}_{m}^{F}(f)=\sum_{k\geqslant 0}\binom{k}{m,n}_{\!\!F}\mathcal{S}_{k}^{F}(f).$ (3.25)
For the additive group law, (3.25) implies $\bigl(\mathcal{S}_{1}^{F_{a}}\bigr)^{\circ n}=n!\cdot\mathcal{S}^{F_{a}}_{n}$, which shows the importance of considering higher hyperderivatives in the presence of torsion. Equation (3.25) follows from the associativity of $F$.
@@ 3.8 $\boldsymbol{F}$-hyperderivatives
$\mathcal{S}_{j}^{F}(z^{m})=\sum_{i\in\mathbb{Z}}\binom{m}{i,j}_{\!\!F}z^{i}.$
From this, we see that (3.25) implies
@@
For all $n\geqslant 1$, we have
$\operatorname{Res}^{F}_{z=0}\mathcal{S}_{n}^{F}f(z){\rm d}z=0,$ (3.26)
$\operatorname{Res}^{F}_{z=0}\mathcal{S}_{n}^{F}f(z)\cdot g(z){\rm d}z=-\sum_{j=1}^{n}\operatorname{Res}_{z=0}^{F}\bigl[\mathcal{S}_{n-j}^{F}f(z)\cdot\mathcal{S}_{j}^{F}g(z)\bigr]{\rm d}z.$ (3.27)
@@
$\operatorname{Res}^{F}_{z=0}i_{z,\underline{w}}f(F(z,w)){\rm d}z=\operatorname{Res}^{F}_{z=0}f(z){\rm d}z.$
We will apply the substitution rule (3.20) to the power series h(z)=F(z,w)wh(z)=F(z,w)-w with coefficients in the ring $R\llbracket w\rrbracket$. This corresponds to viewing $w$ as a fixed constant. Putting ${\rm d}w=0$ in (2.5) gives $p_{F}(F(z,w))F^{1,0}(z,w)=p_{F}(z)$. Clearly, $h(0)=0$ and $h^{\prime}(0)=F^{1,0}(0,w)=p_{F}(w)^{-1}$ is a unit, so the hypotheses for (3.20) are satisfied. Applying this and unravelling the definition gives
$\operatorname{Res}^{F}_{z=0}i_{z,\underline{w}}f(F(z,w)){\rm d}z$ $=\operatorname{Res}_{z=0}i_{z,\underline{w}}f(F(z,w))p_{F}(z){\rm d}z$
$=\operatorname{Res}_{z=0}i_{z,\underline{w}}f(F(z,w))p_{F}(F(z,w))F^{1,0}(z,w){\rm d}z$
$=\operatorname{Res}_{z=0}i_{z,\underline{w}}f(h(z)+w)p_{F}(h(z)+w)h^{\prime}(z){\rm d}z$
$=\operatorname{Res}_{v=0}i_{v,\underline{w}}f(v+w)p_{F}(v+w){\rm d}v.$
Let
$g(z)=f(z)p_{F}(z)=\sum_{n\in\mathbb{Z}}a_{n}z^{n}\in R(\!(z)\!).$
It remains to prove that
$\operatorname{Res}_{v=0}i_{v,\underline{w}}g(v+w){\rm d}v=\operatorname{Res}_{z=0}g(z){\rm d}z,$
which is a straight-forward computation
$\operatorname{Res}_{v=0}i_{v,\underline{w}}g(v+w){\rm d}v=\operatorname{Res}_{v=0}\sum_{n,k\in\mathbb{Z}}a_{n}\binom{n}{k}v^{n-k}w^{k}{\rm d}v=a_{-1},$
@@
For all $f\in R\big\llbracket x_{0}^{\pm 1},x_{1}^{\pm 1},x_{2}^{\pm 1}\big\rrbracket$ such that the substitutions $f(x_{1}-x_{2},x_{1},x_{2})$, $f(x_{0},x_{1},x_{1}-x_{0})$ converge in the algebraic sense,
$\operatorname{Res}_{x_{1}=0}\operatorname{Res}_{x_{2}=0}i_{x_{1},x_{2}}f(x_{1}-x_{2},x_{1},x_{2}){\rm d}x_{2}{\rm d}x_{1}$
$\qquad-\operatorname{Res}_{x_{2}=0}\operatorname{Res}_{x_{1}=0}i_{x_{2},x_{1}}f(x_{1}-x_{2},x_{1},x_{2}){\rm d}x_{1}{\rm d}x_{2}$
$\phantom{\qquad-}{}=\operatorname{Res}_{x_{1}=0}\operatorname{Res}_{x_{0}=0}i_{x_{1},x_{0}}f(x_{0},x_{1},x_{1}-x_{0}){\rm d}x_{0}{\rm d}x_{1}.$ (3.28)
@@
$\operatorname{Res}^{F}_{z_{1}=0}\operatorname{Res}^{F}_{z_{2}=0}i_{z_{1},z_{2}}f(F(z_{1},\iota(z_{2})),z_{1},z_{2}){\rm d}z_{2}{\rm d}z_{1}$
$\qquad-\operatorname{Res}^{F}_{z_{2}=0}\operatorname{Res}^{F}_{z_{1}=0}i_{z_{2},z_{1}}f(F(z_{1},\iota(z_{2})),z_{1},z_{2}){\rm d}z_{1}{\rm d}z_{2}$
$\phantom{\qquad-}{}=\operatorname{Res}^{F}_{z_{1}=0}\operatorname{Res}^{F}_{z_{0}=0}i_{z_{1},z_{0}}f(z_{0},z_{1},F(z_{1},\iota(z_{0}))){\rm d}z_{0}{\rm d}z_{1}.$ (3.29)
@@
$(-1)^{c-1}\binom{a}{1-c}-(-1)^{a+b-1}\binom{a}{1-b}=(-1)^{a-1}\binom{c}{1-a}$
@@
It suffices to prove (3.29) for the universal formal group law, where we may assume (2.8). In this case, we can use (3.21) to reduce (3.29) to (3.28) for the function $f(x_{0},x_{1},x_{2})$, where $\phi(z_{i})=x_{i}$, $i=0,1,2$. ∎
4 Vertex $\boldsymbol{F}$-algebras and their Lie algebras
@@ 4.1 Axioms
A vertex $F$-algebra over a formal group law $F(z,w)\in R\llbracket z,w\rrbracket$ consists of data $(V,\mathbbm{1},\mathcal{S},Y)$ as follows:
  • an $R$-module $V$ of states,
  • a vacuum vector $\mathbbm{1}\in V$,
@@
    $\mathcal{S}(z)\colon V\longrightarrow V\llbracket z\rrbracket,\qquad\mathcal{S}(z)a=\sum_{n\geqslant 0}\mathcal{S}^{(n)}(a)z^{n},$ (4.1)
@@
    $V\otimes_{R}V\longrightarrow V(\!(z)\!),\qquad a\otimes b\longmapsto Y(a,z)b,$
    $Y(a,z)b=\sum_{n\in\mathbb{Z}}a_{(n)}(b)z^{-n-1},\qquad a_{(n)}(b)=0\qquad\text{for}\ n\gg 0.$ (4.2)
@@
  1. $(1)$
    Vacuum & creation: $Y(a,z)\mathbbm{1}$ is holomorphic for all $a\in V$ and
    $Y(a,z)\mathbbm{1}|_{z=0}=a,\qquad Y(\mathbbm{1},z)=\operatorname{id}_{V}.$ (4.3)
  2. $(2)$
    $F$-translation covariance: for all $a,b\in V$, we have
    $Y(\mathcal{S}(w)(a),z)b=i_{z,w}Y(a,F(z,w))b,\qquad\mathcal{S}(z)\mathbbm{1}=\mathbbm{1}.$ (4.4)
    In (4.4), we have used the substitution (3.13). Moreover, $\mathcal{S}(z)\circ\mathcal{S}(w)=\mathcal{S}(F(z,w))$, $\mathcal{S}(0)=\operatorname{id}_{V}$.
  3. $(3)$
    Weak $F$-associativity: for all $a,b,c\in V$, there exists $N\geqslant 0$ with
@@ 3. $(3)$
    using the substitution $v\mapsto F(z,w)$ from (3.11) on the right-hand side.
  4. $(4)$
@@
    $Y(a,z)b=\mathcal{S}(z)\circ Y(b,\iota(z))a.$ (4.6)
Putting $b=\mathbbm{1}$ into (4.6) and using (4.3) gives
$Y(a,z)\mathbbm{1}=\mathcal{S}(z)a.$ (4.7)
@@
The axioms are slightly redundant. For example, (4.4) implies (4.5) for $b=\mathbbm{1}$. On the other hand, we cannot deduce (4.4) from (4.5) unless we know a priori that both sides of (4.4) are meromorphic in $z$, $w$. Hence (4.4) encodes a ‘meromorphicity’ property.
Remark 4.3.
For an ordinary vertex algebra, one usually works with the translation operator $D$, see [11]. The shift operator then is $\mathcal{S}(z)=\exp(zD)$.
Remark 4.4.
@@
The author thanks an anonymous referee for pointing out that, in Huang’s geometric interpretation of ordinary vertex algebras, conformal invariance plays a crucial role. In the genus one setting, an invariant construction requires not only the vertex algebra itself but also its modules and intertwining operators, whose matrix coefficients need not be meromorphic. Full modular invariance is presently established only in the $C_{2}$-cofinite (rational) case due to Huang [7], and more recently in certain $C_{1}$-cofinite settings in work of Creutzig–McRae–Yang [3]. It therefore remains unclear whether a theory of vertex $F$-algebras over an elliptic curve could satisfy such invariance properties, although a formulation with weaker invariance (topological or birational) might still be meaningful.